OpenStudy (ledah):

A parallelogram has a height of 3 m and a base of 6 m. What is its area?

Parth (parthkohli):

\(Area(Parallelogram) = base \times height\)

Parth (parthkohli):

Need more help?

OpenStudy (ledah):

Yeah the easy question seem hard to me because I've got a bad headache

OpenStudy (ledah):

seems*

OpenStudy (ledah):

nah I can take it from there

OpenStudy (anonymous):

Area = Base X Height A=3*6 A=18m^2